# Trilinos: Enabling Scientific Computing Across Diverse Hardware Architectures at Scale

### v1

@@ 4.1. Nonlinear Solvers: NOX, LOCA
NOX (Nonlinear Object-oriented Solutions) provides robust and efficient algorithms for solving systems of nonlinear equations. NOX implements a number of Newton-based globalization techniques including line search (Pawlowski et al., 2006b), trust region (Pawlowski et al., 2006b, 2008), and homotopy algorithms (Coffey et al., 2003). Additionally, it provides lower- and higher-order models including Broyden, Anderson acceleration (Walker and Ni, 2011), and tensor methods (Bader, 2005). The algorithms have been designed for large-scale parallel inexact linear solvers using Krylov methods and support Jacobian-free Newton-Krylov variants. The library interacts with application codes through the Thyra model evaluator interface. NOX provides linear algebra abstractions for applications to use custom implementations with the algorithms. The library additionally contains abstractions for solvers, directions, and line searches that allow users to further customize the algorithms. NOX provides various stopping criteria, including absolute, relative, and weighted root mean square norms, as well as stagnation and NaN detection. Applications can build custom stopping criteria within a tree-based logical structure and provide additional criteria through the StatusTest abstraction. NOX also provides capabilities for continuation and stability analysis through the subpackage LOCA.
@@ 4.2. Numerical Optimization: ROL
Rapid Optimization Library (ROL) (rol, [n. d.]; Javeed et al., 2022) is the Trilinos package for numerical optimization. ROL brings an extensive collection of state-of-the-art optimization algorithms to virtually any application. Its programming interface supports any computational hardware, including heterogeneous many-core systems with digital and analog accelerators. ROL has been used with great success for optimal control, optimal design, inverse problems, image processing, and mesh optimization. Application areas include geophysics, climate science (Perego, 2022), structural dynamics (Aquino et al., 2019; Bunting et al., 2021), fluid dynamics (Antil et al., 2023b), electromagnetics, quantum computing, hypersonics, and geospatial imaging (Kouri and Shields, 2014).
@@
  • Vector abstractions and matrix-free interface for universal applicability: Similar to other Trilinos packages that implement abstract numerical algorithms, ROL’s design is centered around an abstract linear algebra interface, through the ROL::Vector class, which enables the use of any ROL algorithm with any data type, such as the C++ std::vector, MPI-parallel data structures (e.g., Epetra, Tpetra, PETSc, HYPRE vectors), GPU-supporting data structures (e.g., Kokkos, ArrayFire (Yalamanchili et al., 2015)), etc. The abstract linear algebra interface further enables the definition of custom vector dot products, which are critical for performance in large-scale applications. Finally, all ROL algorithms are matrix-free and rely on user-defined applications of linear and nonlinear operators. ROL’s vector class design is inspired by the Hilbert Class Library (Gockenbach and Symes, 1996), the Rice Vector Library (Padula et al., 2009), and the RTOp framework (Bartlett et al., 2004), while ROL’s operator interfaces for simulation-based optimization, known as the SimOpt middleware, are inspired by (Heinkenschloss and Vicente, 1999).
@@
  • Modern algorithms for unconstrained and constrained optimization: ROL features a large collection of well-established algorithms for smooth optimization as well as several novel algorithms developed by the ROL team. ROL categorizes optimization problems into four types based on the presence of certain constraints: Type U (unconstrained), Type B (bound constraints), Type E (equality constraints), and Type G (general constraints). Additionally, each problem type can include linear constraints, which ROL can reduce or eliminate algorithmically. The novel algorithms include augmented Lagrangian methods for infinite-dimensional problems with general constraints (Antil et al., 2023a) and matrix-free trust-region methods for convex-constrained optimization (Kouri, 2022).
@@
  • Easy-to-use methods for stochastic and risk-aware optimization: ROL contains specialized interfaces and algorithms for stochastic optimization. Optimization problems with stochastic and uncertain problem data pose challenges in problem formulation and efficient numerical solution methods. ROL’s interface for stochastic optimization enables the use of numerous modern risk measures, probabilistic functions, and robust problem formulations. To approximate risk-averse problems (Shapiro et al., 2021), ROL includes sample-based approaches such as sample average approximation, stochastic approximation, and adaptive sparse-grid quadrature (Kouri et al., 2013, 2014). ROL also includes specialized algorithms such as progressive hedging (Rockafellar and Wets, 1991) and the primal-dual risk minimization algorithm (Kouri and Surowiec, 2022).
@@
  • Fast and robust algorithms for nonsmooth optimization: ROL includes specialized algorithms to minimize the sum of a smooth function and a nonsmooth, convex function. This class of problems is ubiquitous in computational science. For example, nonsmooth sparsifying regularizers such as the $\ell^{1}$-norm are common in data science and machine learning. These algorithms can be used to solve convex constrained optimization problems. ROL’s methods for this class of problems are built on the proximity operator and include proximal gradient and Newton methods (Beck, 2017; Kanzow and Lechner, 2021; Ochs et al., 2014). As discussed in the subsequent bullet, ROL can also incorporate inexact values and gradients for the smooth objective function term using trust regions (Baraldi and Kouri, 2023). ROL also include an implementation of the trust bundle method (Schramm and Zowe, 1992) for solving general nonconvex, nonsmooth optimization problems.
@@
  • PDE-OPT application development kit for PDE-constrained optimization: To build and optimize models based on partial differential equations (PDEs), ROL offers an application development kit called PDE-OPT. PDE-OPT is based on a modular design with three layers: local finite element computations, which encode the physical equations governing the optimization problem; global computations, which utilize Tpetra data structures; and an interface to ROL, enabling efficient optimization. To solve PDE-constrained optimization problems, users only need to implement the local finite element layer of PDE-OPT (i.e., the evaluations of weak forms on a mesh element); PDE-OPT automates all remaining steps.
@@ 4.3. Automatic Differentiation: Sacado
Sacado (Phipps and Gay, 2011; Phipps and Pawlowski, 2012; Phipps et al., 2008) provides forward and reverse-mode operator-overloading-based automatic differentiation (AD) tools within Trilinos. Sacado’s forward AD tools have been integrated with Kokkos and have demonstrated good performance on GPU architectures (Phipps et al., 2022). Sacado, along with its Kokkos integration, provides high-performance derivative capabilities to numerous Office of Science and NNSA (National Nuclear Security Administration) extreme scale applications, including: Albany for solid mechanics and land ice modeling (Salinger et al., 2016; Hoffman et al., 2018); Charon for semiconductor device modeling (Musson et al., 2020) and multiphase chemically reacting flows (Musson et al., 2009); Drekar for computational fluid dynamics (CFD) (Sondak et al., 2021; Shadid et al., 2016b), magnetohydrodynamics (Shadid et al., 2016a), and plasma physics (Crockatt et al., 2022; Miller et al., 2019); Xyce for electronic circuit simulation (Baker et al., 2012; Keiter et al., 2018); and SPARC for hypersonic fluid flows (Kieweg et al., [n. d.]).
@@ 4.4. Uncertainty Quantification: Stokhos
Stokhos’ implementation of the intrusive stochastic Galerkin uncertainty quantification method allows stochastic projections to be computed, such as polynomial chaos and generalized polynomial chaos expansions. The implementation includes interfaces for forming the linear and/or nonlinear systems that follow from the stochastic Galerkin projection, as well as linear solver methods that can exploit the block structure of these systems. GPU performance is targeted by using Kokkos and by commuting the layout of the Galerkin operator to be outer-spatial and inner-stochastic (Phipps et al., 2014). The implementation has been used in (Constantine et al., 2014) to efficiently propagate uncertainty in multiphysics systems by reducing the full system with a nonlinear elimination method.
The embedded ensemble propagation consists of propagating a subset of samples gathered into a so-called ensemble through the forward simulation at once. It builds on (Pawlowski et al., 2012a) for automating embedded analysis capabilities. Stokhos defines an ensemble type: a SIMD data type that is able to store the values of input, output, and state variables for every sample of an ensemble. This type can then be used within the Tpetra solver stack as a template argument for the scalar type. This approach allows to save computation time in four ways: the sample-independent data and computation can be reused for every sample of an ensemble, the memory access pattern is improved, the operations on the ensemble type can be vectorized efficiently, and the message passing costs are reduced by sending fewer but larger messages. However, the approach requires solvers and BLAS functions to be aware of the extra dimension associated to the ensemble; for example, a GMRES for ensemble types (Liegeois et al., 2020) needs to monitor the convergence of the individual samples in order to decide when to stop based on the union of the information.
@@ 4.5. Nonlinear Analysis Tools: Piro
Piro (Salinger and Phipps, 2010) provides a simple and unified interface to nonlinear analysis tools in Trilinos, exposing capabilities for computing sensitivities, performing parameter continuation and bifurcation analysis, and solving constrained optimization problems. In particular, Piro implements main driver classes for:
  • Linear/nonlinear solvers and sensitivity analysis for transient and nontransient problems. As an example, this capability can be used to compute the discrete solution of a partial differential equation depending on a parameter and the sensitivity of a quantity of interest with respect to this parameter. Sensitivities can be computed in a forward or adjoint fashion, the latter being preferable in the presence of high-dimensional parameters. This capability relies on NOX for the solution of nonlinear problems and Tempus for transient problems.
@@
  • Parameter continuation and bifurcation analysis: this capability is provided through the package LOCA.
These driver classes share a similar interface based on the Thyra::ModelEvaluator and the Teuchos::ParameterList classes. Applications define the problems to be targeted (e.g., the equations to be solved, the parameters, the quantities of interests) by providing a concrete implementations of the Thyra::ModelEvaluator; see (Pawlowski et al., 2012a, b).
@@ 5. Discretization Tools
This section describes Trilinos packages that provide tools for spatial and temporal discretization of integro-differential equations. Most discretization efforts in Trilinos have been devoted to implementing tools for mesh-based discretizations of partial differential equations (PDEs) with a focus on high-order finite elements. A notable exception is the research package Compadre (Section 5.5), which provides tools for meshless approximation of linear operators that can be used for the discretization of differential equations and for data transfer. Trilinos’ discretization packages have been adopted by many applications addressing a wide range of physics problems, including solid mechanics, earth system modeling, semiconductor devices modeling, and electro-magnetics. These applications have taken different approaches in adopting Trilinos mesh-based discretization tools. The less intrusive approach is the adoption of Intrepid2 tools to perform local finite element assembly. In this case, the application has to manage the global assembly, possibly using the DoFManager provided by Panzer, and FECrs matrix and vector structures provided by Tpetra. A more intrusive approach is to additionally use the Phalanx package for managing dependencies of field evaluations in conjunction with the Thyra Model Evaluator and Sacado automatic differentiation, and possibly Tempus for time integration. This approach is particularly useful when developing solvers for complex multiphysics problems, because it allows easy re-use of computational kernels and automates the computation of Jacobians and sensitivities. The most intrusive approach is to build the application around the Panzer package, which provides all of the above, plus the handling of linear and nonlinear solvers and integrated constrained optimization capabilities. In the following, we neither include a description of the snapshotted packages Sierra ToolKit (STK) and Krino, which provide mesh and level-set tools, nor a description of the Shards package, which provides tools for mesh cell topology.
@@ 5.1. Local Assembly: Intrepid2
[…]
@@
[…]

### v2

@@ 4.1. Nonlinear Solvers: NOX, LOCA
NOX (Nonlinear Object-oriented Solutions) provides robust and efficient algorithms for solving systems of nonlinear equations. NOX implements a number of Newton-based globalization techniques including line search (Pawlowski et al., 2006b), trust region (Pawlowski et al., 2006b, 2008), and homotopy algorithms (Coffey et al., 2003). Additionally, it provides lower- and higher-order models including Broyden, Anderson acceleration (Walker and Ni, 2011), and tensor methods (Bader, 2005). The algorithms have been designed for large-scale parallel inexact linear solvers using Krylov methods and support Jacobian-free Newton-Krylov variants. The library interacts with application codes through the Thyra model evaluator interface. NOX provides linear algebra abstractions for applications to use custom implementations with the algorithms. All algorithms are implemented under base-class abstractions for the solver, direction, and line search objects. Users can provide derived implementations for each of these base classes to insert new algorithms into the library. NOX provides various stopping criteria, including absolute, relative, and weighted root mean square norms, as well as stagnation and NaN detection. Applications can build custom stopping criteria within a tree-based logical structure and provide additional criteria through the StatusTest abstraction. NOX also provides capabilities for continuation and stability analysis through the subpackage LOCA.
@@ 4.2. Numerical Optimization: ROL
Rapid Optimization Library (ROL) (124; A. Javeed, D. P. Kouri, D. Ridzal, and G. Von Winckel (2022)) is a package for numerical optimization. ROL brings an extensive collection of state-of-the-art optimization algorithms to virtually any application. Its programming interface supports modern computational hardware, including heterogeneous multi-core systems. ROL has been used with great success for optimal control, optimal design, inverse problems, image processing, and mesh optimization. Application areas include geophysics, climate science (Perego, 2022), structural dynamics (Aquino et al., 2019; Bunting et al., 2021), fluid dynamics (Antil et al., 2023a), electromagnetics, quantum computing, hypersonics, and geospatial imaging (Kouri and Shields, 2014).
@@
  • Vector abstractions and matrix-free interface for universal applicability: Similar to other Trilinos packages that implement abstract numerical algorithms, ROL’s design is centered around an abstract linear algebra interface, through the ROL::Vector class, which enables the use of any ROL algorithm with any data type, such as the C++ std::vector, MPI-parallel data structures (e.g., Tpetra, PETSc, HYPRE vectors), GPU-supporting data structures (e.g., Kokkos, ArrayFire (Yalamanchili et al., 2015)), etc. In contrast to other vector implementations, ROL’s abstract linear algebra interface also enables the definition of custom vector dot products with duality pairings, which are critical for performance in large-scale optimization with differential equations. Finally, all ROL algorithms are matrix-free and rely on user-defined applications of linear and nonlinear operators. ROL’s vector class design is inspired by the Hilbert Class Library (Gockenbach and Symes, 1996), the Rice Vector Library (Padula et al., 2009), and the RTOp framework (Bartlett et al., 2004), while ROL’s operator interfaces for simulation-based optimization, known as the SimOpt middleware, are inspired by (Heinkenschloss and Vicente, 1999).
@@
  • Modern algorithms for unconstrained and constrained optimization: ROL features a large collection of well-established algorithms for smooth optimization as well as several novel algorithms developed by the ROL team. ROL categorizes optimization problems into four types based on the presence of certain constraints: Type U (unconstrained), Type B (bound constraints), Type E (equality constraints), and Type G (general constraints). Additionally, each problem type can include linear constraints, which ROL can reduce or eliminate algorithmically. The novel algorithms include augmented Lagrangian methods for infinite-dimensional problems with general constraints (Antil et al., 2023b) and matrix-free trust-region methods for convex-constrained optimization (Kouri, 2022).
@@
  • Easy-to-use methods for stochastic and risk-aware optimization: ROL contains specialized interfaces and algorithms for stochastic optimization. Optimization problems with stochastic and uncertain problem data pose challenges in problem formulation and efficient numerical solution methods. ROL’s interface for stochastic optimization enables the use of numerous modern risk measures, probabilistic functions, and robust problem formulations. To approximate stochastic problems (Shapiro et al., 2021), ROL includes sample-based approaches such as sample average approximation, stochastic approximation, and adaptive sparse-grid quadrature (Kouri et al., 2013, 2014). ROL also includes specialized algorithms such as progressive hedging (Rockafellar and Wets, 1991) and the primal-dual risk minimization algorithm (Kouri and Surowiec, 2022).
@@
  • Fast and robust algorithms for nonsmooth optimization: ROL includes specialized algorithms to minimize the sum of a smooth function and a nonsmooth, convex function. This class of problems is ubiquitous in computational science. For example, nonsmooth sparsifying regularizers such as the $\ell^{1}$-norm are common in data science and machine learning (Fuhg et al., 2024). These algorithms can be used to solve convex constrained optimization problems. ROL’s methods for this class of problems are built on the proximity operator and include proximal gradient and Newton methods (Beck, 2017; Kanzow and Lechner, 2021; Ochs et al., 2014). As discussed in the subsequent bullet, ROL can also incorporate inexact values and gradients for the smooth objective function term using trust regions (Baraldi and Kouri, 2023). ROL also include an implementation of the trust bundle method (Schramm and Zowe, 1992) for solving general nonconvex, nonsmooth optimization problems.
@@
  • PDE-OPT application development kit for PDE-constrained optimization: To build and optimize models based on PDEs, ROL offers an application development kit for finite element models, called PDE-OPT. PDE-OPT is based on a modular design with three layers: local finite element computations, which encode the physical equations governing the optimization problem; global computations, which utilize Tpetra data structures; and an interface to ROL, enabling efficient optimization. To solve PDE-constrained optimization problems, users only need to implement the local finite element layer of PDE-OPT (i.e., the evaluations of weak forms on a mesh element); PDE-OPT automates all remaining steps.
In 2025, the development of ROL has moved to its own GitHub repository [^13] https://github.com/sandialabs/rol. The ROL code is snapshotted into the Trilinos repository for convenience of Trilinos users.
@@ 4.3. Automatic Differentiation: Sacado
Sacado (Phipps and Gay, 2011; Phipps and Pawlowski, 2012; Phipps et al., 2008) provides forward and reverse-mode operator-overloading-based automatic differentiation (AD) tools within Trilinos. Sacado’s forward AD tools have been integrated with Kokkos and have demonstrated good performance on GPU architectures (Phipps et al., 2022). Sacado, along with its Kokkos integration, provides high-performance derivative capabilities to numerous Office of Science and NNSA (National Nuclear Security Administration) extreme scale applications, including: Albany for solid mechanics and land ice modeling (Salinger et al., 2016; Hoffman et al., 2018); Charon for semiconductor device modeling (Musson et al., 2020) and multiphase chemically reacting flows (Musson et al., 2009); Drekar for computational fluid dynamics (CFD) (Sondak et al., 2021; Shadid et al., 2016b), magnetohydrodynamics (Shadid et al., 2016a), and plasma physics (Crockatt et al., 2022; Miller et al., 2019); Xyce for electronic circuit simulation (Baker et al., 2012; Keiter et al., 2018); and SPARC for hypersonic fluid flows (Kieweg et al., ).
@@ 4.4. Uncertainty Quantification: Stokhos
Stokhos’ implementation of the intrusive stochastic Galerkin uncertainty quantification method allows stochastic projections to be computed, such as polynomial chaos and generalized polynomial chaos expansions. The implementation includes interfaces for forming the linear and/or nonlinear systems that follow from the stochastic Galerkin projection, as well as linear solver methods that can exploit the block structure of these systems. In particular, stochastic Galerkin methods applied to PDEs employ a tensor product discretization between the spatial and stochastic discretizations that generates a two-level sparsity pattern in the resulting stochastic Galerkin operator. This operator is traditionally organized in a block fashion where the outer sparsity pattern is dictated by the stochastic discretization while each block follows the sparsity pattern of the PDE spatial discretization. To improve performance on architectures such as GPUs, Stokhos commutes this layout to obtain an outer sparsity pattern dictated by the spatial discretization where each block follows the stochastic discretization and leverages Kokkos hierarchical parallelism to implement calculations corresponding to the stochastic discretization in a scalable and efficient manner (Phipps et al., 2014). The implementation has been used in (Constantine et al., 2014) to efficiently propagate uncertainty in multiphysics systems by reducing the full system with a nonlinear elimination method.
The embedded ensemble propagation consists of propagating a subset of samples gathered into a so-called ensemble through the forward simulation at once. It builds on (Pawlowski et al., 2012b) for automating embedded analysis capabilities. Stokhos defines an ensemble type: a data type that is able to store the values of input, output, and state variables for every sample of an ensemble and can take advantage of SIMD (single instruction, multiple data) operations. This type can then be used within the Tpetra solver stack as an argument for its scalar type template parameter, allowing instantiation of Tpetra data structures on the ensemble type to propagate ensembles through supported linear solvers and preconditioners. This approach results in reduced computation time in four ways: the sample-independent data and computation can be reused for every sample of an ensemble, memory access patterns can be improved, operations on the ensemble type can be vectorized efficiently, and message passing costs can be reduced by sending fewer but larger messages. However, the approach requires solvers and BLAS functions to be aware of the extra dimension associated to the ensemble; for example, GMRES for ensemble types (Liegeois et al., 2020) needs to monitor the convergence of the individual samples in order to decide when to terminate based on whether every sample within the ensemble has satisfied the chosen stopping criteria.
@@ 4.5. Nonlinear Analysis Tools: Piro
Piro (Salinger and Phipps, 2010) provides a simple and unified interface to nonlinear analysis tools in Trilinos, exposing capabilities for computing sensitivities, performing parameter continuation and bifurcation analysis, and solving constrained optimization problems. Piro wraps many of the common functionalities of Trilinos packages such as Belos, NOX, LOCA, Tempus and ROL, into a simple interface. In particular, Piro implements main driver classes for:
  • Linear/nonlinear solvers and sensitivity analysis for transient and nontransient problems. As an example, this capability can be used to compute the discrete solution of a partial differential equation depending on a parameter and the sensitivity of a quantity of interest with respect to this parameter. Sensitivities can be computed in a forward or adjoint fashion, the latter being preferable in the presence of high-dimensional parameters. To this end, Piro performs forward and adjoint solves, relying on Belos, NOX and Tempus for the solution of linear, nonlinear and transient problems.
@@
  • Parameter continuation and bifurcation analysis. The core capabilities for continuation and bifurcation analysis are provided through the package LOCA. The Piro driver integrates these functionalities with linear, nonlinear, and time integration solvers, enhancing flexibility and usability.
These driver classes share a similar interface based on the Thyra::ModelEvaluator and the Teuchos::ParameterList classes. Applications define the problems to be targeted (e.g., the equations to be solved, the parameters, the quantities of interests) by providing a concrete implementations of the Thyra::ModelEvaluator; see (Pawlowski et al., 2012b, a).
@@ 5. Discretization Tools
This section describes Trilinos packages that provide tools for spatial and temporal discretization of integro-differential equations. Most discretization efforts in Trilinos have been devoted to implementing tools for mesh-based discretizations of PDEs with a focus on high-order finite elements. A notable exception is the research package Compadre (Section 5.5), which provides tools for meshless approximation of linear operators that can be used for the discretization of differential equations and for data transfer. Trilinos’ discretization packages have been adopted by many applications addressing a wide range of physics problems, including solid mechanics, earth system modeling, semiconductor devices modeling, and electro-magnetics. These applications have taken different approaches in adopting Trilinos mesh-based discretization tools. The less intrusive approach is the adoption of Intrepid2 tools to perform local finite element assembly. In this case, the application has to manage the global assembly, possibly using the DoFManager provided by Panzer, and FECrs matrix and vector structures provided by Tpetra. A more intrusive approach is to additionally use the Phalanx package for managing dependencies of field evaluations in conjunction with the Thyra Model Evaluator and Sacado automatic differentiation, and possibly Tempus for time integration. This approach is particularly useful when developing solvers for complex multiphysics problems, because it allows easy re-use of computational kernels and automates the computation of Jacobians and sensitivities. The most intrusive approach is to build the application around the Panzer package, which provides all of the above, plus the handling of linear and nonlinear solvers and integrated constrained optimization capabilities. In the following, we neither include a description of the snapshotted packages Sierra ToolKit (STK) and Krino, which provide mesh and level-set tools, nor a description of the Shards package, which provides tools for mesh cell topology.
@@ 5.1. Local Assembly: Intrepid2
Intrepid2 provides interoperable tools for compatible discretizations of PDEs; it is a performance-portable re-implementation and extension of the legacy Intrepid package (Bochev et al., 2012). Intrepid2 mainly focuses on local assembly of continuous and discontinuous finite elements. It also provides limited capabilities for finite volume discretization. Intrepid2 works on batches of elements (cells), and provides tools to efficiently compute discretized linear functionals (e.g., right-hand-side vectors) and differential operators (e.g., stiffness matrices) at the element level. Intrepid2 implements compatible finite element spaces of various polynomial orders for $H({\rm grad})$, $H({\rm curl})$, $H({\rm div})$, and $L^{2}$ function spaces on triangles, quadrilaterals, tetrahedrons, hexahedrons, wedges, and pyramids. It provides both Lagrangian basis functions and hierarchical basis functions (Fuentes et al., 2015), and it implements performance optimizations (e.g., sum factorizations) exploiting the underlying structure of the problem (e.g., tensor-product elements or other symmetries). The degrees of freedom of $H({\rm div})$ and $H({\rm curl})$ finite elements as well as high-order $H({\rm grad})$ finite elements depend on the global orientation of edges and faces and Intrepid2 provides orientation tools for matching the degrees of freedom on shared edges and faces. It also provides interpolation-based projection tools for projecting functions in $H({\rm grad})$, $H({\rm curl})$, $H({\rm div})$, and $L^{2}$ to the respective discrete spaces. Intrepid2 implements these capabilities through the following classes:
@@
  • IntegrationTools: This class provides integration methods that can take advantage of tensor product structures in basis values, providing mechanisms for performance-portable, sum-factorized assembly across $H({\rm grad})$, $H({\rm curl})$, $H({\rm div})$, and $L^{2}$ function spaces. In the future, we plan to provide similar interfaces to support matrix-free discretizations.